# Normalizing Flow-Assisted Nested Sampling on Type-II Seesaw Model

### v1

@@ III.2.1 Set-up
To solve the first problem, we should both run and train the ML algorithms iteratively. The second problem can be solved by using ML predictions to calculate the real likelihoods by running HEP programs, albeit for a much smaller number of points. The solution to the third problem is generating the parameters and theoretical values of observables (or corresponding likelihoods) simultaneously. This can be achieved by training a generative model to draw samples from the joint probability distribution $P(\hat{\theta},\hat{O})$ of the parameters $\hat{\theta}$ and the target observables $\hat{O}$. It can then ‘generate’ random instances (outcomes) of these together [117, 118]. Any competent generative algorithm can be trained on a vector dataset containing parameters and observables (or $\chi^{2}$s or $\mathcal{L}$s).
@@
III.2.2 Classification
After collecting an adequate number of points from the target parameter space that pass the theoretical constraints deriving from vacuum stability and perturbative unitarity [81], we label them either ‘$1$’ or ‘00’, depending on whether the corresponding spectrum file is generated or not. To set up an optimal binary classifier, we first use a variable-structure ‘Self-normalizing Neural Network (SNN)’ as shown on the left of Fig. 2, in a way very similar to that explained in ref.s [22, 119]. As SNNs can only be used with standardized inputs, a feature extractor function is coupled to the network as an encoder. The hyper-parameters varied are: the number of modules ($n_{m}$), the number of neurons in each module ($n_{w}$), and the $\alpha$-dropout probability ($p_{\alpha}$) of the specialized self-normalizing rectilinear unit (SELU). The hyper-parameters were varied over these ranges (initial value, final value, step size):
@@ III.2.2 Classification
The best combination was found to be $(n_{m},n_{w},p_{\alpha})=(4,200,0.007)$, with an accuracy of $99.38\%$ with an ADAM optimizer [120], which increases to $99.47\%$ with a subsequent training of the same network with a stochastic gradient descent (SGD) optimizer.
@@
III.2.3 Real NVP as a Generative ML model
Several different Generative Algorithms are at our disposal at present, including but not limited to: Generative Adversarial Networks (GANs) [121], Variational Auto-encoders (VAEs) [122], and Flow-based generative models [123, 124], also called Normalizing Flows (NFs). Neither GANs nor VAEs can learn the probability density of the training data and as will be apparent in this section, it is an integral part of our ML speed-up, although we will not use the learned PDFs directly as our result. We have used a special type of NF, called Real NVP (RNVP; Real-valued Non-Volume Preserving transformations) [125] in our work.
@@ III.2.3 Real NVP as a Generative ML model
“In each simple bijection, part of the input vector is updated using a function which is simple to invert, but which depends on the remainder of the input vector in a complex way.” [125]
@@ III.3.1 Initialization and NS
After the initial inspection, the starting dataset is prepared with $>63000$ vectors of 13 components, for which the total $\chi^{2}$ is not very large ($\leq 1335$). This dataset is used to create a pool of points (labeled ‘Collected points’ in figure 1) to train the first iteration of the RNVP generative algorithm, as well as a starting Nested Sample (by using the dataset as a representation of the actual parameter-space). At this starting point, the saved sample only has points sampled from the low-likelihood region and the tolerance is $\approx 50$.
III.3.2 RNVP and Training
@@
The generative performance of the RNVP depends on the distribution of the training samples. A generator, trained only once in the beginning, has a considerably low probability of generating high-likelihood points, as neither the position nor the volume of the high-likelihood region is a priori known. The computational expense too, for it to be representative enough, is unknown and probably as large as a naive NS. Even in the rare and ideal case of a large and sufficiently representative starting dataset, using it whole (or an iid smaller sample) to train the generator for every iteration would be as prohibitive as using the entire prior for the NS (see sec. III.3.1).
As the nature of the constrained prior changes non-trivially in exploration problems, the best bet is to iteratively train the generator on a dataset sampled from the constrained prior within the shrinking bounding region. As training is time-consuming, and its predictions do not change considerably for the inner high-likelihood region, it is profitable to train the generator onece after several NS iterations. Because we save all the truth vectors from the HEP packages in the pooled dataset, the number of samples saved in that pool is a good measure to decide when to train again. We have chosen to train the RNVP network after a group of $6000$ points are dumped in the pool every time.
@@
Figure 5: Two-dimensional marginal distribution of the posteriors. The contours of the same color are from the same posterior. The solid and dot-dashed contours correspond to $68\%$ respectively, and the similar $95\%$ credible interval contours are dashed and dotted. Red and blue colors enclose all valid points and those valid points allowed by HiggsBounds respectively.
III.3.3 Dataset Selection
The pool (labeled ‘Collected points’ in Fig. 1) contains samples from the whole prior. As was discussed in the previous section, predictions from the distribution learned from this become prohibitive. The predictive speed of the generator depends on the way we choose the training dataset from it (labeled ‘Database Selection’ in Fig. 1). We have run many instances of the ML-NS algorithm to find the optimal method applicable to any BSM model, of which we exhibit the main three competing options here:
@@ III.3.3 Dataset Selection
  • To use the iso-likelihood contours with $\mathcal{L}$ slightly lower than the lowest of the live points, instead of the bounding region, for dataset-selection. When the constrained prior volume is small, it faithfully represents the prior, while being much faster in generating new, high-likelihood points. As we will see, it produces a high bias in the generated sample if used from the beginning, as a result of not correctly representing the prior. We will call it $RNVP-F$ from now on.
@@
Fig. 4 contains the $\Delta Z$ vs. time curve of three such runs: the orange (dashed) curve represents the $RNVP-S$ one, the green (dot-dashed) curve the $RNVP-F$ and the blue (solid) curve the $RNVP-SF$. The $RNVP-S$ works quite fast at the start but once $\Delta Z$ falls below $10$, it slows down by a lot, adding $3-4$ points to the NS-sample per hour. We truncated the plot here at $150$ hrs, but, in reality, we ran it for $>700$ hrs, with a final tolerance of $\sim 4.09$, when it was finally stopped. While being extremely slow, it was successfully converging to the correct posterior. The $RNVP-F$ run, on the other hand, reached the termination tolerance quite fast and at a steady rate. However, the posterior it generated was wrong (it peaked at a different place than the concentration of the live points, i.e., the high-likelihood region). The first part of the $RNVP-SF$ is identical to $RNVP-S$, but here, when the sample rate (number of NS-samples correctly added between two consecutive trainings) falls below 50 (at $\sim 56$ hrs) as shown in the plot on the right of Fig. 4, we start the $RNVP-F$ procedure. This resulted in a considerable speed-up in the algorithm, without changing the nature of the target posterior perceptibly. The final tolerance reached by both $RNVP-F$ and $RNVP-SF$ is $<0.001$.
Figure 7: All points allowed by all theoretical and experimental constraints up to $95\%$ CL: (a) in the doubly and singly charged Higgs mass plane (Top-left), (b) in the singly charged and neutral heavy Higgs mass plane (top-right) (c) The mixing angle between the CP-even Higgs states as a function of the triplet VEV (Bottom-left), and (d) The cross-section $\sigma\times B(H^{\pm\pm}H^{\mp\mp}\to W^{\pm}W^{\pm}W^{\mp}W^{\mp})$ in fb unit as a function of $m_{H^{\pm\pm}}$ along with the $1\sigma$ and $2\sigma$ expected limits [127] (Bottom-right).
IV Results
In this section, we discuss the results obtained through our analysis and highlight their salient features and importance. We have only explored a relatively high $v_{T}$ region ($\sim 10^{-4}-2.5$). With this choice, the quartic couplings and the $\mu_{1}$ parameter were explored to the fullest in this analysis. As mentioned earlier, the objective here was to locate the high-likelihood regions given the present experimental data on the SM-like Higgs boson mass, signal strengths, and the $\rho$-parameter along with the oblique parameters. After locating the favored parameter space, we have shown the impact of direct scalar (neutral and charged) search results on the obtained parameter space using HiggsBounds. We do the same for the doubly charged scalar with the latest LHC results. All regions of the PS shown in this section obey the vacuum stability and perturbative unitarity conditions.
IV.1 Posteriors of the parameter space
@@
The top-left figure shows the distribution in the $\mu_{1}$ vs. $v_{T}$ plane. The red contours in the figure indicate the boundaries of the $68\%$ (dot-dashed) and $95\%$ (dotted) credible intervals of the valid sample from the posterior. The blue curves indicate similar boundaries with all points from the sample allowed by HiggsBounds (solid: $68\%$, dashed: $95\%$). The shaded contours with a blue color gradient paint a more detailed picture of the high-likelihood region (the darker the shade, the higher the likelihood). Thus, the darkest shaded region indicates the most favored area of the parameter space in this parameter plane. Evidently, any $\mu_{1}$ values beyond $\sim 70$ GeV is disfavored by the collider data at $95\%$ interval. The triplet VEV is rather tightly constrained with no valid points outside the $0.05-2.6$ GeV range subjected to the value of $\mu_{1}$. The highest likelihood region of $v_{T}$ however is restricted within $\sim$ 1.2-1.4 GeV. This is mostly due to the $\rho$-parameter.
The figure on the top right panel of Fig. 5 shows the $2D$ marginal posterior in $\lambda$ vs. $v_{T}$ plane. The quartic coupling $\lambda$ is tightly constrained mostly from the SM-like Higgs mass measurement while the triplet VEV on this parameter plane is even more constrained than in the previous figure. This apparent slight change in the credible intervals is a numerical artifact of the $2D$ marginalization of a seven-dimensional space. The HiggsBounds constraints on the obtained sample flattens the marginal modes a little and thus, the recalculated $68\%$ and $95\%$ credible intervals slightly overshoot the corresponding unconstrained (red) contours.
@@
IV.2 Observable Space
IV.2.1 Physical Masses and properties of the scalars
With a proper idea of the favored parameter space at our disposal, let us discuss their impact on the physical scalar masses. As was pointed out in Sec. II, in addition to one singly- and one doubly-charged scalar, the Type-II seesaw model has one CP-odd and two CP-even neutral scalars.
The top panel of Fig. 7 shows the favored parameter space in different BSM Higgs mass planes. In the top-left figure, the $y$ and $x$ axes show the mass difference between the singly and doubly charged Higgs states and the doubly charged Higgs mass, respectively. Only the points allowed by all theoretical and phenomenological constraints at $95\%$ CL are shown here. It is evident from the figure that most $m_{H^{\pm\pm}}$ values are concentrated in the $\sim 450-800$ GeV range. The mass differences give a clear idea about the corresponding $m_{H^{\pm}}$. Electroweak precision study with the oblique parameters leads to a limit on the mass difference $|m_{H^{\pm\pm}}-m_{H^{\pm}}|\lesssim 40$ GeV [52]. We obtained the same window for the mass difference without applying this constraint a priori, although the points seem to bunch up within the $0-20$ GeV mass gap. The figure on the top-right compares the heavy CP-even Higgs masses ($m_{H}$) with that of the singly charged Higgs masses. The masses are evidently quite close, lying within a mass gap of $\sim 20$ GeV throughout the favored PS. This is expected since the mass squared differences $m^{2}_{H^{\pm}}-m^{2}_{H}$ and $m^{2}_{H^{\pm\pm}}-m^{2}_{H^{\pm}}$ are nearly identical under certain approximations for the range of $v_{T}$ we have obtained here[48]. We have not shown any distribution with the CP-odd Higgs state since it is mass-degenerate with the heavier CP-even state. This is expected since we have not allowed any CP violation in the scalar sector. As is evident from Eq. 7, the mixing angle ($\alpha$) between the two CP-even Higgs states depends on all seven input parameters. The bottom-left panel of Fig. 7 shows how $\sin\alpha$ varies as a function of $v_{T}$. Understandably, $\sin\alpha$ has to be small owing to the present-day Higgs data and the figure clearly shows that it can be at most $\sim 10^{-3}$. All points in this figure are allowed by all theoretical and phenomenological constraints at the $95\%$ CL. Note, that any smaller mixing angle, though allowed in principle, will not fit the SM-like Higgs data within $95\%$ CL.
@@ IV.2.1 Physical Masses and properties of the scalars
Note that, all scalars apart from the doubly charged Higgs state have already been subjected to direct search constraints through HiggsBounds. To comment on the impact of direct search results of doubly charged Higgs states on the favored parameter space obtained here, we have to first talk about possible decays of $H^{\pm\pm}$. Since we are working with a large $v_{T}$ scenario, the Yukawa couplings have to be extremely tiny to be consistent with neutrino data, as indicated by Eq.12. All LNV decays are, therefore, highly suppressed over all of the favored parameter space. As a result, the only possible two-body decay mode of this scalar is $H^{\pm\pm}\to W^{\pm}W^{\pm}$ which dominates over all possible three-body decay modes. The other possible two-body decays e.g., into $H^{\pm}H^{\pm}$ or $H^{\pm}W^{\pm}$ are not kinematically allowed due to the small mass gap of $m_{H^{\pm\pm}}$ and $m_{H^{\pm}}$. We have checked the consistency of the favored parameter space against the latest collider search results of doubly charged Higgs; produced via vector boson fusion (VBF)[128, 129, 130] or pair-produced via photon or $Z$-boson mediated diagrams [127]. They decay into $W^{\pm}W^{\pm}$ in both cases. The cross-section was computed using MadGraph [131, 132]. We find that for the VBF production, the whole favored PS is quite safe from the existing results with the cross-section $\sigma_{\rm VBF}(H^{\pm\pm})\times(H^{\pm\pm}\to W^{\pm}W^{\pm})$ lying at least one order of magnitude below the experimental limit. The doubly charged Higgs can also be pair-produced at the LHC with both of them decaying into $W$-bosons [127]. The experimental collaboration studied the experimental observation in the context of the Type-II seesaw model and showed that $m_{H^{\pm\pm}}\lesssim 350$ GeV is excluded from this search at 13 TeV center-of-mass energy with the luminosity at $139{\rm fb}^{-1}$ luminosity. Almost all the points allowed by theoretical and phenomenological considerations up to $95\%$ CLs produce $m_{H^{\pm\pm}}\gtrsim 400$ GeV with a sparse region in between $350-400$ GeV. We show the comparison with the experimental result on the bottom-right of Fig. 7. We have shown the $1\sigma$ and $2\sigma$ expected limit and the $95\%$ expected upper limit for $\sigma\times BR(pp\to H^{\pm\pm}H^{\mp\mp}\to W^{\pm}W^{\pm}W^{\mp}W^{\mp})$ along with the theoretical prediction obtained for Type-II seesaw model [127]. The red line represents the cross-section calculated at next-to-leading-order by the experimental collaboration and the dotted black line shows our leading-order calculation multiplied by a k-factor of 1.32. We can safely conclude from this figure that the favored parameter space is yet to be explored by the collider search of $H^{\pm\pm}$.
IV.2.2 Yukawa matrix from neutrino oscillation data

### v2

@@ 3.2.1 Set-up
To solve the first problem, we should both run and train the ML algorithms iteratively. The second problem can be solved by using ML predictions to calculate the real likelihoods by running HEP programs, albeit for a much smaller number of points. The solution to the third problem is generating the parameters and theoretical values of observables (or corresponding likelihoods) simultaneously. This can be achieved by training a generative model to draw samples from the joint probability distribution $P(\hat{\theta},\hat{O})$ of the parameters $\hat{\theta}$ and the target observables $\hat{O}$. It can then ‘generate’ random instances (outcomes) of these together mitchell1997machine ; 10.5555/2980539.2980648 . Any competent generative algorithm can be trained on a vector dataset containing parameters and observables (or $\chi^{2}$s or $\mathcal{L}$s).
@@
3.2.2 Classification
After collecting an adequate number of points from the target parameter space that pass the theoretical constraints deriving from vacuum stability and perturbative unitarity Primulando:2019evb , we label them either ‘$1$’ or ‘00’, depending on whether the corresponding spectrum file is generated or not. To set up an optimal binary classifier, we first use a variable-structure ‘Self-normalizing Neural Network (SNN)’ as shown on the left of Fig. 2, in a way very similar to that explained in ref.s Baruah:2024gwy ; Bhattacharya:2020vme . As SNNs can only be used with standardized inputs, a feature extractor function is coupled to the network as an encoder. The hyper-parameters varied are: the number of modules ($n_{m}$), the number of neurons in each module ($n_{w}$), and the $\alpha$-dropout probability ($p_{\alpha}$) of the specialized self-normalizing rectilinear unit (SELU). The hyper-parameters were varied over these ranges (initial value, final value, step size):
@@ 3.2.2 Classification
The best combination was found to be $(n_{m},n_{w},p_{\alpha})=(4,200,0.007)$, with an accuracy of $99.38\%$ with an ADAM optimizer kingma2014adam , which increases to $99.47\%$ with a subsequent training of the same network with a stochastic gradient descent (SGD) optimizer.
@@
3.2.3 Real NVP as a Generative ML model
Several different Generative Algorithms are at our disposal at present, including but not limited to: Generative Adversarial Networks (GANs) goodfellow2014generative , Variational Auto-encoders (VAEs) kingma2013auto , and Flow-based generative models papamakarios2021normalizing ; kobyzev2020normalizing , also called Normalizing Flows (NFs). Neither GANs nor VAEs can learn the probability density of the training data and as will be apparent in this section, it is an integral part of our ML speed-up, although we will not use the learned PDFs directly as our result. We have used a special type of NF, called Real NVP (RNVP; Real-valued Non-Volume Preserving transformations) dinh2016density in our work.
@@ 3.2.3 Real NVP as a Generative ML model
“In each simple bijection, part of the input vector is updated using a function which is simple to invert, but which depends on the remainder of the input vector in a complex way.” dinh2016density
@@ 3.3.1 Initialization and NS
After the initial inspection, the starting dataset is prepared with $>63000$ vectors of 13 components, for which the total $\chi^{2}$ is not very large ($\leq 1335$). This dataset is used to create a pool of points (labelled ‘Collected points’ in figure 1) to train the first iteration of the RNVP generative algorithm, as well as a starting Nested Sample (by using the dataset as a representation of the actual parameter-space). At this starting point, the saved sample only has points sampled from the low-likelihood region and the tolerance is $\approx 50$.
3.3.2 RNVP and Training
@@
The generative performance of the RNVP depends on the distribution of the training samples. A generator, trained only once in the beginning, has a considerably low probability of generating high-likelihood points, as neither the position nor the volume of the high-likelihood region is a priori known. The computational expense too, for it to be representative enough, is unknown and probably as large as a naive NS. Even in the rare and ideal case of a large and sufficiently representative starting dataset, using it whole (or an iid smaller sample) to train the generator for every iteration would be as prohibitive as using the entire prior for the NS (see sec. 3.3.1).
As the nature of the constrained prior changes non-trivially in exploration problems, the best bet is to iteratively train the generator on a dataset sampled from the constrained prior within the shrinking bounding region. As training is time-consuming, and its predictions do not change considerably for the inner high-likelihood region, it is profitable to train the generator once after several NS iterations. Because we save all the truth vectors from the HEP packages in the pooled dataset, the number of samples saved in that pool is a good measure to decide when to train again. We have chosen to train the RNVP network after a group of $6000$ points is dumped in the pool every time.
@@
Figure 5: Two-dimensional marginal distribution of the posteriors. The contours of the same colour are from the same posterior. The solid and dot-dashed contours correspond to $68\%$ respectively, and the similar $95\%$ credible interval contours are dashed and dotted. Red and blue colours enclose all valid points and those valid points allowed by HiggsBounds respectively.
3.3.3 Dataset Selection
The pool (labelled ‘Collected points’ in Fig. 1) contains samples from the whole prior. As was discussed in the previous section, predictions from the distribution learned from this become prohibitive. The predictive speed of the generator depends on the way we choose the training dataset from it (labelled ‘Database Selection’ in Fig. 1). We have run many instances of the ML-NS algorithm to find the optimal method applicable to any BSM model, of which we exhibit the main three competing options here:
@@ 3.3.3 Dataset Selection
  • To use the iso-likelihood contours with $\mathcal{L}$ slightly lower than the lowest of the live points, instead of the bounding region, for dataset selection. When the constrained prior volume is small, it faithfully represents the prior, while being much faster in generating new, high-likelihood points. As we will see, it produces a high bias in the generated sample if used from the beginning, as a result of not correctly representing the prior. We will call it $RNVP-F$ from now on.
@@
Fig. 4 contains the $\Delta Z$ vs. time curve of three such runs: the orange (dashed) curve represents the $RNVP-S$ one, the green (dot-dashed) curve the $RNVP-F$ and the blue (solid) curve the $RNVP-SF$. The $RNVP-S$ works quite fast at the start but once $\Delta Z$ falls below $10$, it slows down by a lot, adding $3-4$ points to the NS-sample per hour. We truncated the plot here at $150$ hrs., but, in reality, we ran it for $>700$ hrs., with a final tolerance of $\sim 4.09$, when it was finally stopped. While being extremely slow, it was successfully converging to the correct posterior. The $RNVP-F$ run, on the other hand, reached the termination tolerance quite fast and at a steady rate. However, the posterior it generated was wrong (it peaked at a different place than the concentration of the live points, i.e., the high-likelihood region). The first part of the $RNVP-SF$ is identical to $RNVP-S$, but here, when the sample rate (number of NS-samples correctly added between two consecutive trainings) falls below 50 (at $\sim 56$ hrs.) as shown in the plot on the right of Fig. 4, we start the $RNVP-F$ procedure. This resulted in a considerable speed-up in the algorithm, without changing the nature of the target posterior perceptibly. The final tolerance reached by both $RNVP-F$ and $RNVP-SF$ is $<0.001$.
Figure 7: All points allowed by all theoretical and experimental constraints up to $95\%$ CL: (a) in the doubly and singly charged Higgs mass plane (Top-left), (b) in the singly charged and neutral heavy Higgs mass plane (top-right) (c) The mixing angle between the CP-even Higgs states as a function of the triplet VEV (Bottom-left), and (d) The cross-section $\sigma\times B(H^{\pm\pm}H^{\mp\mp}\to W^{\pm}W^{\pm}W^{\mp}W^{\mp})$ in $fb$ unit as a function of $m_{H^{\pm\pm}}$ along with the $1\sigma$ and $2\sigma$ expected limits ATLAS:2021jol (Bottom-right).
4 Results
In this section, we discuss the results obtained through our analysis and highlight their salient features and importance. We have only explored a relatively high $v_{T}$ region ($\sim 10^{-4}-2.5$). With this choice, the quartic couplings and the $\mu_{1}$ parameter were explored to the fullest in this analysis. As mentioned earlier, the objective here was to locate the high-likelihood regions given the present experimental data on the SM-like Higgs boson mass, signal strengths, and the $\rho$-parameter along with the oblique parameters. After locating the favoured parameter space, we have shown the impact of direct scalar (neutral and charged) search results on the obtained parameter space using HiggsBounds. We do the same for the doubly charged scalar with the latest LHC results. All regions of the PS shown in this section obey the vacuum stability and perturbative unitarity conditions.
4.1 Posteriors of the parameter space
@@
The top-left figure shows the distribution in the $\mu_{1}$ vs. $v_{T}$ plane. The red contours in the figure indicate the boundaries of the $68\%$ (dot-dashed) and $95\%$ (dotted) credible intervals of the valid sample from the posterior. The blue curves indicate similar boundaries with all points from the sample allowed by HiggsBounds (solid: $68\%$, dashed: $95\%$). The shaded contours with a blue colour gradient paint a more detailed picture of the high-likelihood region (the darker the shade, the higher the likelihood). Thus, the darkest shaded region indicates the most favoured area of the parameter space in this parameter plane. Evidently, any $\mu_{1}$ values beyond $\sim 70$ GeV is disfavored by the collider data at $95\%$ interval. The triplet VEV is rather tightly constrained with no valid points outside the $0.05-2.6$ GeV range subjected to the value of $\mu_{1}$. The highest likelihood region of $v_{T}$ however is restricted within $\sim$ 1.2-1.4 GeV. This is mostly due to the $\rho$-parameter.
The figure on the top right panel of Fig. 5 shows the $2D$ marginal posterior in $\lambda$ vs. $v_{T}$ plane. The quartic coupling $\lambda$ is tightly constrained mostly from the SM-like Higgs mass measurement while the triplet VEV on this parameter plane is even more constrained than in the previous figure. This apparent slight change in the credible intervals is a numerical artefact of the $2D$ marginalization of a seven-dimensional space. The HiggsBounds constraints on the obtained sample flatten the marginal modes a little and thus, the recalculated $68\%$ and $95\%$ credible intervals slightly overshoot the corresponding unconstrained (red) contours.
@@
4.2 Observable Space
4.2.1 Physical Masses and properties of the scalars
With a proper idea of the favoured parameter space at our disposal, let us discuss their impact on the physical scalar masses. As was pointed out in Sec. 2, in addition to one singly- and one doubly-charged scalar, the Type-II seesaw model has one CP-odd and two CP-even neutral scalars.
The top panel of Fig. 7 shows the favoured parameter space in different BSM Higgs mass planes. In the top-left figure, the $y$ and $x$ axes show the mass difference between the singly and doubly charged Higgs states and the doubly charged Higgs mass, respectively. Only the points allowed by all theoretical and phenomenological constraints at $95\%$ CL are shown here. It is evident from the figure that most $m_{H^{\pm\pm}}$ values are concentrated in the $\sim 450-800$ GeV range. The mass differences give a clear idea about the corresponding $m_{H^{\pm}}$. Electroweak precision study with the oblique parameters leads to a limit on the mass difference $|m_{H^{\pm\pm}}-m_{H^{\pm}}|\lesssim 40$ GeV Chun:2012jw . We obtained the same window for the mass difference without applying this constraint a priori, although the points seem to bunch up within the $0-20$ GeV mass gap. The figure on the top-right compares the heavy CP-even Higgs masses ($m_{H}$) with that of the singly charged Higgs masses. The masses are evidently quite close, lying within a mass gap of $\sim 20$ GeV throughout the favoured PS. This is expected since the mass squared differences $m^{2}_{H^{\pm}}-m^{2}_{H}$ and $m^{2}_{H^{\pm\pm}}-m^{2}_{H^{\pm}}$ are nearly identical under certain approximations for the range of $v_{T}$ we have obtained hereMelfo:2011nx . We have not shown any distribution with the CP-odd Higgs state since it is mass-degenerate with the heavier CP-even state. This is expected since we have not allowed any CP violation in the scalar sector. As is evident from Eq. 7, the mixing angle ($\alpha$) between the two CP-even Higgs states depends on all seven input parameters. The bottom-left panel of Fig. 7 shows how $\sin\alpha$ varies as a function of $v_{T}$. Understandably, $\sin\alpha$ has to be small owing to the present-day Higgs data and the figure clearly shows that it can be at most $\sim 10^{-3}$. All points in this figure are allowed by all theoretical and phenomenological constraints at the $95\%$ CL. Note, that any smaller mixing angle, though allowed in principle, will not fit the SM-like Higgs data within $95\%$ CL.
@@ 4.2.1 Physical Masses and properties of the scalars
Note that, all scalars apart from the doubly charged Higgs state have already been subjected to direct search constraints through HiggsBounds. To comment on the impact of direct search results of doubly charged Higgs states on the favoured parameter space obtained here, we have to first talk about possible decays of $H^{\pm\pm}$. Since we are working with a large $v_{T}$ scenario, the Yukawa couplings have to be extremely tiny to be consistent with neutrino data, as indicated by Eq.12. All LNV decays are, therefore, highly suppressed over all of the favoured parameter space. As a result, the only possible two-body decay mode of this scalar is $H^{\pm\pm}\to W^{\pm}W^{\pm}$ which dominates over all possible three-body decay modes. The other possible two-body decays e.g., into $H^{\pm}H^{\pm}$ or $H^{\pm}W^{\pm}$ are not kinematically allowed due to the small mass gap of $m_{H^{\pm\pm}}$ and $m_{H^{\pm}}$. We have checked the consistency of the favoured parameter space against the latest collider search results of doubly charged Higgs; produced via vector boson fusion (VBF)ATLAS:2024txt ; ATLAS:2023dbw ; CMS:2021wlt or pair-produced via photon or $Z$-boson mediated diagrams ATLAS:2021jol . They decay into $W^{\pm}W^{\pm}$ in both cases. The cross-section was computed using MadGraph Alwall:2014hca ; Frederix:2018nkq . We find that for the VBF production, the whole favoured PS is quite safe from the existing results with the cross-section $\sigma_{\rm VBF}(H^{\pm\pm})\times(H^{\pm\pm}\to W^{\pm}W^{\pm})$ lying at least one order of magnitude below the experimental limit. The doubly charged Higgs can also be pair-produced at the LHC with both of them decaying into $W$-bosons ATLAS:2021jol . The experimental collaboration studied the experimental observation in the context of the Type-II seesaw model and showed that $m_{H^{\pm\pm}}\lesssim 350$ GeV is excluded from this search at 13 TeV centre-of-mass energy with the luminosity at $139{\rm fb}^{-1}$ luminosity. Almost all the points allowed by theoretical and phenomenological considerations up to $95\%$ CLs produce $m_{H^{\pm\pm}}\gtrsim 400$ GeV with a sparse region in between $350-400$ GeV. We show the comparison with the experimental result on the bottom-right of Fig. 7. We have shown the $1\sigma$ and $2\sigma$ expected limit and the $95\%$ expected upper limit for $\sigma\times BR(pp\to H^{\pm\pm}H^{\mp\mp}\to W^{\pm}W^{\pm}W^{\mp}W^{\mp})$ along with the theoretical prediction obtained for Type-II seesaw model ATLAS:2021jol . The red line represents the cross-section calculated at next-to-leading-order by the experimental collaboration and the dotted black line shows our leading-order calculation multiplied by a k-factor of 1.32. We can safely conclude from this figure that the favoured parameter space is yet to be explored by the collider search of $H^{\pm\pm}$.
4.2.2 Yukawa matrix from neutrino oscillation data